OpenStudy (anonymous):

How would you find the slope and y-intercept of y = -2/3x -12

OpenStudy (amistre64):

y = slope x + intercept

OpenStudy (anonymous):

Y=mx+b where m is the slope and b is the intercept

OpenStudy (radar):

With the above knowledge, you can solve this by inspection.

OpenStudy (anonymous):

does anyone have an actual example

OpenStudy (anonymous):

It is driving me absolutely mad

OpenStudy (radar):

Use the problem as an example. What is m the slope. You know it is the coefficient of x!

OpenStudy (amistre64):

there is a number in the place for "slope"; what number is that?

OpenStudy (amistre64):

there is a number in place of the intercept, what number is that?

OpenStudy (anonymous):

-4, -10 does this make sense?

OpenStudy (amistre64):

not quite, at least not with the problem you provided

OpenStudy (anonymous):

Dont I just plug in my own numbers?

OpenStudy (amistre64):

y = -2/3 x + -12 y = slope x + intercept

OpenStudy (amistre64):

this question is designed to get you to become familiar with the form of the equation; not any particular points on it

OpenStudy (amistre64):

slope intercept uses the slope and the intercept as part of the equation itself

OpenStudy (anonymous):

oh my gosh, okay, thank you! I wil keep working until I get it

OpenStudy (amistre64):

good luck :)

OpenStudy (radar):

When the line equation is in slope intercept form, as is the one you are working with, the coefficient of x is the slope, the constant is the intercept. They are there in your equation, just pick them out.

OpenStudy (anonymous):

So what radar is saying is that you don't need to do anything. The answer is right in front of you. Many times in math we complicate things too much and many times the answer is right in front of us.

OpenStudy (anonymous):

So in this case m is the slope. so what is M? Then what is B since B is the y intercept?

OpenStudy (anonymous):

Maybe this will help y= (m)x + (b)

OpenStudy (anonymous):

oh my gosh, thank you so much, my daughter is here helping me and it is right in front of my face! That is so true, Mathdude!!

OpenStudy (anonymous):

Thank you so much, Radar...it is slowly....very very slowly coming together!

OpenStudy (anonymous):

Sorry, I have been out of school for so long, I have forgotten!